This is a multi-step equation. I need the steps and the answer. What is r: 3r/4-4/5=7/10
These are all fractions.

Answer by richwmiller(17219)

3r/4-4/5=7/10
multiply through by 20
5*3r-4*4=7*2
15r-16=14
Can you take it from here?